Banks in New Transylvania have a desired reserve ratio of 10 percent of deposits and no excess reserves. The currency drain ratio is 50 percent of deposits. Now suppose that the central bank increases the monetary base by $1,200 billion.

  1. How much do the banks lend in the first round of the money creation process?
  2. How much of the initial amount loaned flows back to the banking system as new
  3. deposits?
  4. How much of the initial amount loaned does not return to the banks but is held
  5. as currency?
  6. Why does a second round of lending occur?
  7. Calculate money multiplier in this example?
  8. What is the final increase in the quantity of money?
1
Expert's answer
2022-03-22T11:25:46-0400

Solution

i.). In the first round of the money creation process, the banks lend out the same amount as the increase in the monetary base. Therefore, in this case, the banks lend out $1,200 billion and this is also the increase in the new money.

 

ii.). Since the currency drain ratio is 50% of deposits, one-third of the new money will be held as currency and two thirds will be held as new deposits:

New deposits ="\frac{2}{3} \times 1,200 = \$800" billion

 

iii.). Amount held as currency = "\frac{1}{3} \times 1,200 = \$400" billion

 

iv.). Second round of lending occurs in the event that all loans have been collected and the banks have excess reserves.

 

v.). Money Multiplier = "\frac{1}{1 - (1-r) (1-c)}"


Where: r = Reserve ratio = 10%

           c = Currency drain ratio = 50%

MM = "\frac{1}{1 - (1-0.1) (1-0.5)} = \frac{1}{1 - (0.9) (0.5)} = \frac{1}{1 - (0.45} = 1.82"


Money Multiplier = 1.82

 

vi.). Change in quantity of money = Money multiplier x Change in the monetary base.

Change in monetary base = 1,200 billion

Change in quantity of money = "1.82\times1,200" = 2, 184

Change in quantity of money = $2, 184 billion